suppose W is plane 3x +6y-4z=0. a basis for W is?



1
Expert's answer
2021-10-21T12:43:53-0400

Writing z in terms of x and y,


"-4z=-3x-6y"

"z=\frac{3}{4}x+\frac{3}{2}y"


Any vector basis,

"(x,y,z)=(x,y,\frac{3}{4}x+\frac{3}{2}y)"

When "x=1,y=0,z=\frac{3}{4}"

Setting "x=0,y=1,z=\frac{3}{2}"

"(x,y,\frac{3}{4}x+\frac{3}{2}y)"

"=x (1,0,\frac{3}{4})+y(0,1,\frac{3}{2})"


Basis is given by


"\{\begin{bmatrix}
 1\\
 0\\¾
\end{bmatrix},\begin{bmatrix}
 0\\
 1\\3/2
\end{bmatrix}\}"